# YuLan-Mini: An Open Data-efficient Language Model

### v1

@@ 1 Introduction
Figure 1: Performance comparison of YuLan-Mini against other base models, based on the average scores across eight benchmarks: GSM8K, MATH-500, HumanEval, MBPP, MMLU, ARC-Challenge, HellaSwag, and CEval. Floating Point Operations (FLOPs) are estimated using the formula $C=6ND$, as described in Kaplan et al. (2020). The models with a size larger than 3B are plotted in grey.
@@
To demonstrate the effectiveness of our pre-trained base model, we conduct extensive experiments on a variety of benchmarks, and compare it with a few competitive base models from both research and industry. Experimental results show that our base model, YuLan-Mini, can achieve very promising results among these compared models. For instance, it (the 28K version) achieves scores of 37.80 on MATH-500 (four-shot), 64.00 on HumanEval (zero-shot), and 49.10 on MMLU (five-shot). Figure 1 presents a comparison of YuLan-Mini with other industry models [^3] We select eight popular benchmarks to cover math, coding, general, and language capabilities. This figure is primarily intended to illustrate the training efficacy of our model, rather than to serve as an accurate capacity ranking of existing models..
@@ 2.1 Model Architecture
We adopt rotary positional embedding (ROPE) to capture the positional information in our model, since it integrates absolute and relative positioning in an unified way. During the stable training stage, we set the parameter $\theta$ to 10 0001000010\,00010 000, and increase it to 49 0004900049\,00049 000 during the annealing stage to extend the context length to 28 6722867228\,67228 672 (28K) tokens using adjusted base frequency (ABF).
@@ 2.2 Tokenizer
Generally, the vocabulary size should be chosen to balance its effects on the model’s parameter size and efficiency. We adopt the three approaches proposed by Dagan et al. (2024) to balance the compute budget and vocabulary capacity, yielding a final vocabulary size of around 99,000. For simplicity, we reuse the Byte Pair Encoding (BPE) tokenizer of MiniCPM (Hu et al., 2024). Specifically, we truncate the vocabulary by applying the corresponding BPE merge rules to reduce the number of tokens. We also heuristically remove rare domain-specific tokens, while add some reserved tokens in the vocabulary. The statistics of the modified vocabulary and the compression rate are shown at Table 2. The test set for the tokenization experiments is derived from a diverse array of datasets, as detailed in Section 6.3. Overall, our tokenization method achieves a well-balanced compression rate across different domains.
@@ 3.1.2 Theoretical Analysis and Empirical Evidence on Exploding Hidden States
where $\mathbf{Z}$ denotes the scaled attention scores. The base dimensionality $d_{\text{model}}$ of LLMs are often large (e.g., 1,920 in our model). Additionally, the vanilla setup in Hugging Face uses a default initialization standard deviation of 0.02 for the weight matrices. When training a proxy model, this initialization leads to significantly large variance values, exacerbated by the effects of the aforementioned Attention and FFN modules. As a result, the gradient norm becomes very large. A detailed derivation of the attention head can be found in Takase et al. (2023). Common mitigation strategies include initializing the weights with a very small variance, typically inversely proportional to $d_{\text{model}}$, which we will discuss in Section 3.2.
@@ 3.2.1 Scaled Initialization and Scaling Factor
As discussed in Section 3.1.2, employing a carefully designed initialization method along with appropriate scaling factors is a key strategy for addressing training instability. In this regard, we explore the initialization strategies proposed by Megatron-LM (Shoeybi et al., 2020) and BLOOM (Scao et al., 2022). Specifically, we initialize $\mathbf{W}_{2}$ of the FFN layer (Equation (1)) and $\mathbf{W}_{o}$ of the Attention layer (Equation (2)) in accordance with $\mathcal{N}(0,\sigma^{2}/(2n_{\text{layers}}))$, and initialize the remaining parameters according to $\mathcal{N}(0,\sigma^{2})$. Here, we set the initialization standard deviation as $\sigma=\sqrt{2/(5d)}$.
@@ 3.2.2 Maximal Update Parametrization
In the above, we have explored the importance and effectiveness of parameter initialization through theoretical analysis and empirical experiments. However, when we migrate the training configuration selected on the proxy model to the model of the target size, instability still exists.
@@ 3.3.1 Warmup Based Methods
Learning rate warmup involves gradually increasing the learning rate from a small initial value (e.g., 00) to the max learning rate in $T_{\text{LR}}$ steps. Wortsman et al. (2024) suggests that a longer learning rate warmup can reduce sensitivity to the learning rate, as measured by training stability across different learning rates. We empirically verify this conclusion and find increasing $T_{LR}$ indeed enhances training stability. For our final training, we set $T_{\text{LR}}=2433$, which approximately corresponds to $10$ billion tokens of data.
@@ 4.1 Data Collection
Encyclopedia Wikipedia, Baidu-Baike 14.80B
Books CBook, Gutenberg, LoC-PD-Books 52.13B
@@
General Knowledge arXiv, StackExchange, CICG-news 121.87B
Synthetic Data (Pretrain & Ours) Synthesized math related texts (seed: AutoMathText) 8.76B
Synthetic Data (Instruction & Ours) Synthesized CoT (seed: MetaMathQA) 23.52B
@@ 4.3.1 Math Reasoning
We also incorporate reasoning data from formal mathematics, such as formal theorem proving, which has been shown to enhance performance on general mathematical tests like GSM8K and MATH. Specifically, we collect Lean code dataset like DeepSeek-Prover (Xin et al., 2024) and its associated prover states to train the model in generating proof tactics. For the Lean GitHub dataset (Wu et al., 2024), we concatenate together the reasoning steps in each document to form a long reasoning chain. Additionally, inspired by LIME (Wu et al., 2021), we augment the Lean Workbook (Ying et al., 2024) datasets with three reasoning primitives: (1) Deduction: $\text{State}_{\text{before}},\text{Tactic}\rightarrow\text{State}_{\text{after}}$; (2) Abduction: $\text{State}_{\text{after}},\text{Tactic}\rightarrow\text{State}_{\text{before}}$; and (3) Induction: $\text{State}_{\text{before}},\text{State}_{\text{after}}\rightarrow\text{Tactic}$. Rather than directly predicting the next proof tactics, we train the model to predict the previous or next state based on the proof tactics.
@@ 4.3.2 Code Reasoning
To enhance existing programming competition datasets, like those from LeetCode, [^6] https://huggingface.co/datasets/greengerong/leetcode we utilize the in-context learning (ICL) method by leveraging a small number of demonstrations. This approach allows us to generate additional examples, thereby expanding and diversifying the data. By introducing a broader range of challenging programming problems, we enrich the dataset significantly.
@@ 5.1 Optimization Setting
$f(n,N,N_{\text{annealing}})=\left(1-\sqrt{\frac{n-(N-N_{\text{annealing}})}{N_{\text{annealing}}}}\right),$
@@
Following the work by Hu et al. (2024), we estimate the optimal annealing ratio to be 8%, i.e., 80 billion tokens. We maintain the same batch size used during stable training, i.e., 4 million. The learning rate is decreased from $10^{-2}$ to $5.22\times 10^{-5}$ over a span of 18,802 steps. Subsequently, the learning rate is held constant at $5.22\times 10^{-5}$ for the final 772 steps.
@@
Previous research (Chen et al., 2023) has demonstrated that LLMs can hardly process texts exceeding their context windows due to the out-of-distribution rotation angles in RoPE. To achieve the context window extension, increasing the base frequency of RoPE to migrate the OOD rotation angles and continual pre-training has been an effective method (Xiong et al., 2024). Consequently, during the annealing stage, we increase the base frequency of RoPE $\theta$ from 10,000, employed during stable training, to 490,000 and train the model on long texts. This adjustment successfully extends the context length from 4,096 (4K) tokens to 28,672 (28K) tokens.
@@ 6.1.1 Evaluation Benchmarks
  • Language comprehension: We select the widely-used English benchmarks MMLU (Hendrycks et al., 2021a) and LAMBADA (Kazemi et al., 2023), along with the Chinese benchmarks CMMLU (Li et al., 2024a) and CEval (Huang et al., 2023), RACE  (Lai et al., 2017), to evaluate the bilingual comprehension capabilities of the LLM. These benchmarks span various domains, such as history, science, and culture.
@@ 6.1.3 Implementation Details
  • Chain-of-Thought (CoT): For GSM8K and MATH, we follow previous work Qwen-Team (2024) that uses CoT prompting to facilitate the LLM to perform step-by-step reasoning. Considering the potential performance variance caused by CoT prompts, we utilize both the short ones provided by the original dataset and the long ones generated by kimi-k0-math. For each model, we evaluate the performance using both prompt types, and select the one yielding the higher score as the result.
@@
  • Maximum length: For GSM8K and MATH, since CoT prompting may result in longer outputs, we set the maximum generation length to 596 for short context models and 2,048 for long context models. For HumanEval and MBPP, we set the maximum generation length to 400. For other generative tasks, we set it to 128 for efficiency.
@@ Appendix A Definition of Variables
$n_{\text{heads}}$ The num of model’s layers
$n_{\text{kv\_heads}}$ The num of model’s layers
$d_{\text{model,base}}$ Proxy (base) model’s layer width, hyper-parameter
$d_{\text{model}}$ Dimension of each layer
$d_{\text{head}}$ Dimension of each attention head
$d_{\text{ffn}}$ The feed-forward network’s hidden size
$\sigma_{\text{base}}$ Base initialization standard deviation for each matrix
$\eta_{\text{base}}$ The base learning rate : Maximum in training schedule
$m_{\text{width}}$ A multiplicative or fully-connected weights tensor
@@ B.1 Experiment Setup
The detailed architecture settings are listed in Table 9. A 0.2B proxy model is utilized for general exploratory experiments, whereas a 0.4B model is employed for muP experiments. The latter proxy model has the same number of layers as the final model.
@@ Appendix E Detailed Data Composition by Training Phases
The specific composition of data for each course stage is shown in Table 13, where black represents English web pages and general content, green represents Chinese, blue represents code, and red represents mathematics. The first 10 billion tokens of Phase 1 are used during the warm-up stage. The following 30 billion tokens of Phase 1, along with Phase 2 through Phase 25, are employed in the stable training stage. Phases 26 and 27 are the annealing stage.

### v2

@@ 1 Introduction
Figure 1: Performance comparison of YuLan-Mini against other base models, based on the average scores across eight benchmarks: GSM8K, MATH-500, HumanEval, MBPP, MMLU, ARC-Challenge, HellaSwag, and CEval. Floating Point Operations (FLOPs) are estimated using the scaling law formula $C=6ND$ proposed by Kaplan et al. (2020), where $N$ is the model size and $D$ is the size of the dataset. The models with a size larger than 3B are plotted in gray.
@@
To demonstrate the effectiveness of our pre-trained base model, we conduct extensive experiments on a variety of benchmarks, and compare it with a few competitive base models from both research and industry. Experimental results show that our base model, YuLan-Mini, can achieve very promising results among these compared models. For instance, it (the 28K version) achieves scores of 37.80 on MATH-500 (four-shot), 64.00 on HumanEval (zero-shot), and 49.10 on MMLU (five-shot). Figure 1 presents a comparison of YuLan-Mini with other industry models. [^3] We select eight popular benchmarks to cover math, coding, general, and language capabilities. This figure is primarily intended to illustrate the training efficacy of our model, rather than to serve as an accurate capacity ranking of existing models.
@@ 2.1 Model Architecture
We adopt rotary positional embedding (ROPE) to capture the positional information in our model, since it integrates absolute and relative positioning in an unified way. During the stable training stage, we set the parameter $\theta$ to 10,000, and increase it to 49 0004900049\,00049 000 during the annealing stage to extend the context length to 28,672 (28K) tokens using adjusted base frequency (ABF).
@@ 2.2 Tokenizer
Generally, the vocabulary size should be chosen to balance its effects on the model’s parameter size and efficiency. We adopt the three approaches proposed by Dagan et al. (2024) to balance the compute budget and vocabulary capacity, yielding a final vocabulary size of around 99,000. For simplicity, we reuse the Byte Pair Encoding (BPE) tokenizer of MiniCPM (Hu et al., 2024). Specifically, we truncate the vocabulary by applying the corresponding BPE merge rules to reduce the number of tokens. We also heuristically remove rare domain-specific tokens, while add some reserved tokens in the vocabulary. The statistics of the modified vocabulary and the compression rate are shown at Table 2. The test set for the tokenization experiments is sourced from a diverse array of datasets, as detailed in Section 6.3. Overall, our tokenization method achieves a well-balanced compression rate across different domains.
@@ 3.1.2 Theoretical Analysis and Empirical Evidence on Exploding Hidden States
where $\mathbf{Z}$ denotes the scaled attention scores. The base dimensionality $d_{\text{model}}$ of LLMs are often large (e.g., 1,920 in our model). Additionally, the vanilla setup in Hugging Face uses a default initialization standard deviation of 0.02 for the weight matrices. When training the proxy model, this initialization leads to significantly large variance values, exacerbated by the effects of the aforementioned Attention and FFN modules. As a result, the gradient norm becomes very large. A detailed derivation of the attention head can be found in Takase et al. (2023). Common mitigation strategies include initializing the weights with a very small variance, typically inversely proportional to $d_{\text{model}}$, which we will discuss in Section 3.2.
@@ 3.2.1 Scaled Initialization and Scaling Factor
As discussed in Section 3.1.2, employing a carefully designed initialization method along with appropriate scaling factors is a key strategy for addressing training instability. In this regard, we explore the initialization strategies proposed by Megatron-LM (Shoeybi et al., 2020) and BLOOM (Scao et al., 2022). Specifically, we initialize $\mathbf{W}_{2}$ of the FFN layer (Equation (1)) and $\mathbf{W}_{o}$ of the Attention layer (Equation (2)) in accordance with $\mathcal{N}(0,\sigma_{\text{base}}^{2}/(2n_{\text{layers}}))$, and initialize the remaining parameters according to $\mathcal{N}(0,\sigma_{\text{base}}^{2})$. Here, we set the initialization standard deviation as $\sigma_{\text{base}}=\sqrt{2/(5d)}$.
@@ 3.2.2 Maximal Update Parametrization
In the above, we have explored the importance and effectiveness of parameter initialization through theoretical analysis and empirical experiments. However, when we migrate the training configuration selected on the 0.05B proxy model to the model of the target size, instability still exists.
@@ 3.3.1 Warmup Based Methods
Learning rate warmup involves gradually increasing the learning rate from a small initial value (e.g., 00) to the max learning rate in $T_{\text{LR}}$ steps. Wortsman et al. (2024) suggests that a longer learning rate warmup can reduce sensitivity to the learning rate, as measured by training stability across different learning rates. We empirically verify this conclusion and find increasing $T_{LR}$ indeed enhances training stability. For our final training, we set $T_{\text{LR}}=\,$2,433, which approximately corresponds to $10$ billion tokens of data.
@@ 4.1 Data Collection
General Knowledge arXiv, StackExchange, English News 121.87B
Books CBook, Gutenberg, LoC-PD-Books 52.13B
Encyclopedia Wikipedia, Baidu-Baike 14.80B
@@
Synthetic Pretrain Data (Ours) Synthetic document (seed: AutoMathText, LeetCode) 8.76B
Synthetic Instruction (Ours) Reasoning (seed: MetaMathQA, DeepMind Math, …) 23.52B
@@ 4.3.1 Math Reasoning
We also incorporate reasoning data from formal mathematics, such as formal theorem proving, which has been shown to enhance performance on general mathematical tests like GSM8K and MATH. Specifically, we collect Lean tactic (e.g., intro, simp) dataset like DeepSeek-Prover (Xin et al., 2024) and its associated prover states to train the model in generating proof tactics. For the Lean GitHub dataset (Wu et al., 2024), we concatenate together the reasoning steps in each document to form a long reasoning chain. Additionally, inspired by LIME (Wu et al., 2021), we augment the Lean Workbook (Ying et al., 2024) datasets with three reasoning primitives: (1) Deduction: $\text{State}_{\text{before}},\text{Tactic}\rightarrow\text{State}_{\text{after}}$; (2) Abduction: $\text{State}_{\text{after}},\text{Tactic}\rightarrow\text{State}_{\text{before}}$; and (3) Induction: $\text{State}_{\text{before}},\text{State}_{\text{after}}\rightarrow\text{Tactic}$. Rather than directly predicting the next proof tactics, we train the model to predict the previous or next state based on the proof tactics.
@@ 4.3.2 Code Reasoning
To enhance existing programming competition datasets, like those from LeetCode [^6] https://huggingface.co/datasets/greengerong/leetcode, we utilize the in-context learning (ICL) method by leveraging a small number of demonstrations. This approach allows us to generate additional examples, thereby expanding and diversifying the data. By introducing a broader range of challenging programming problems, we enrich the dataset significantly.
@@ 5.1 Optimization Setting
$f(n;N,N_{\text{annealing}})=\left(1-\sqrt{\frac{n-(N-N_{\text{annealing}})}{N_{\text{annealing}}}}\right),$
@@
Following the work by Hu et al. (2024), we estimate the optimal annealing ratio to be 8%, i.e., 80 billion tokens. We maintain the same batch size used during stable training, i.e., 4 million tokens. The learning rate is decreased from $10^{-2}$ to $5.22\times 10^{-5}$ over a span of 18,802 steps. Subsequently, the learning rate is held constant at $5.22\times 10^{-5}$ for the final 772 steps.
@@
Previous research (Chen et al., 2023) has demonstrated that LLMs can hardly process texts exceeding their context windows due to the out-of-distribution (OOD) rotation angles in RoPE. To achieve the context window extension, increasing the base frequency of RoPE to migrate the OOD rotation angles and continual pre-training has been an effective method (Xiong et al., 2024). Consequently, during the annealing stage, we increase the base frequency of RoPE $\theta$ from 10,000, employed during stable training, to 490,000 and train the model on long texts. This adjustment successfully extends the context length from 4,096 (4K) tokens to 28,672 (28K) tokens.
@@ 6.1.1 Evaluation Benchmarks
  • Language comprehension: We select the widely-used English benchmarks MMLU (Hendrycks et al., 2021a), LAMBADA (Kazemi et al., 2023) and RACE (Lai et al., 2017), along with the Chinese benchmarks CMMLU (Li et al., 2024a) and CEval (Huang et al., 2023), to evaluate the bilingual comprehension capabilities of the LLM. These benchmarks span various domains, such as history, science, and culture.
@@ 6.1.3 Implementation Details
  • Chain-of-Thought (CoT): For GSM8K and MATH, we follow previous work (Qwen-Team, 2024) that uses CoT prompting to facilitate the LLM to perform step-by-step reasoning. Considering the potential performance variance caused by CoT prompts, we utilize both the short ones provided by the original dataset and the long ones generated by kimi-k0-math. For each model, we evaluate the performance using both prompt types, and select the one yielding the higher score as the result.
@@
  • Maximum length: For GSM8K and MATH, since CoT prompting may result in longer outputs, we set the maximum generation length to 596 for short context (i.e., 4K) models and 2,048 for long context models. For HumanEval and MBPP, we set the maximum generation length to 400. For other generative tasks, we set it to 128 for efficiency.
@@ Appendix A Definition of Variables
$n_{\text{heads}}$ The num of model’s attention heads
$n_{\text{kv\_heads}}$ The num of model’s kv-heads used in GQA
$d_{\text{model}}$ Model dimension, i.e., hidden size
$d_{\text{head}}$ Dimension of attention head
$d_{\text{ffn}}$ The hidden size of feed-forward network
$\sigma_{\text{base}}$ Initialization standard deviation for each matrix
$\eta_{\text{base}}$ Learning rate, i.e., max learning rate
$d_{\text{model\_proxy}}$ $d_{\text{model}}$ for proxy model, i.e., the 0.05B model
$m_{\text{width}}$ Width scaling factor in $\mu$P, i.e., $d_{\text{model}}/d_{\text{model\_proxy}}$
@@ B.1 Experiment Setup
The detailed architecture settings are listed in Table 9. The 0.2B proxy model is utilized for general exploratory experiments, whereas the 0.05B and 0.4B model are employed for $\mu$P experiments. The latter proxy model has the same number of layers as the final model.
@@ Appendix E Detailed Data Composition by Training Phases
The specific composition of data for each course stage is shown in Table 13, where black represents English web pages and general content, green represents Chinese, blue represents code, and red represents mathematics. The first 10 billion tokens of Phase 1 are used during the warm-up stage. The next 30 billion tokens of Phase 1, along with Phases 2 through 25, are employed in the stable training stage. Phases 26 and 27 constitute the annealing stage.